a crafts store bought 46 1/3 ounces of potpouri.The shop divide the potpouri into 1 1/3 ounce bags and sell each bag for $2.50.
A. How many 1 1/3 ounce bags of potpouri can the shop make?
B. If all the bags are sold, how much money will the shop take in?

Answer by mananth(16949)

Total potpouri = 46 1/3 ounces
small packing = 1 1/3 ounces
cost $2.50 each
...
46 1/3 = 139/3
..
1 1/3 = 4/3
Number of small bags = 139/3 / 4/3
=139/3 * 3/4
=139/4
=34.75 bags
34 bags.
..
each cost $2.50
34*2.50 = $85.00